OpenStudy (anonymous):

Plz help!! : A system of equations is shown below: 2x = 5y + 4 3x - 2y = -16 What is the solution to this system of equations?

OpenStudy (anonymous):

A. (-8, -4) B. (8, 4) C. (-4, -8) D. (4, 8)

OpenStudy (bee_see):

you just plug in x and y from your choices and see which one works.

OpenStudy (bee_see):

if you didn't have the choices, then make the equations look the same and get rid of a variable.

OpenStudy (texaschic101):

2x = 5y + 4 2x - 5y = 4 now we have : 2x - 5y = 4 ---->(3) 3x - 2y = -16 --->(-2) ------------- 6x - 15y = 12 (result of multiplying by 3) -6x + 4y = 32 (result of multiplying by -2) ------------- add -11y = 44 y = - 4 2x = 5y + 4 2x = 5(-4) + 4 2x = -20 + 4 2x = -16 x = -8 solution is : (-8,-4)

OpenStudy (texaschic101):

do you have any questions at all ?

OpenStudy (anonymous):

no thx :)